OpenStudy (anonymous):

The surface area of a cube is 100 square centimeters. Estimate to the nearest tenth of a centimeter the side length of the cube using the formula S = 6s2, where S represents surface area and s represents side length. Make sure to show all work and label your answer.

OpenStudy (amistre64):

whats giving you difficulties?

OpenStudy (anonymous):

all of it i dont understand algebra very much lol

OpenStudy (amistre64):

start by filling in the formula, what do we get?

OpenStudy (anonymous):

would we get 16 = 6 *16 ^2 im not sure if thats right or not

OpenStudy (amistre64):

no the formula is: S = 6s^2 they say S stands for Surface Area, and that Surface area is equal to 100 100 = 6s^2 is how we would start to find "s" we need to start unwrapping it ....

OpenStudy (amistre64):

would you agree that 6/6 = 1? and that multipling by 1 is redundant?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

since we want to keep things equal and balanced ... whatever we do to one side of the "+" we need to do to the other; so ... i would suggest we start by dividing each side by 6 100/6 = (6/6) s^2 100/6 = 1 * s^2 100/6 = s^2

OpenStudy (amistre64):

now the question is, how do we undo that ^2?

OpenStudy (amistre64):

ugh .... each side of the "="

OpenStudy (anonymous):

lol. um would we break it down so its like this s*s?

OpenStudy (amistre64):

thats is something we could do, but doesnt really lead us anywhere. the undoing of a square is a sqaure root 3^2 = 9 sqrt(9) = sqrt(3^2) = 3 in general: sqrt(n^2) = n

OpenStudy (amistre64):

100/6 = s^2 , lets sqrt each side sqrt(100/6) = sqrt(s^2) sqrt(100/6) = s